Suppose in the milk store, milk sales per day are represented by a random variable X which follows a normal distribution with.

What is the probability that milk sold per day will be between (a) 185 and 245 gallons? And (b) more than 265 gallons?

Mean: 225 gallons

SD: 35 gallons

Please provide a clear explanation with answers.
Answer #1

a) P(185 < X < 245)

= P(-1.14 < Z < 0.57)

= P(Z < 0.57) - P(Z < -1.14)

= 0.7157 - 0.1271

= 0.5886

b) P(X > 265)

= P(Z > 1.14)

= 1 - P(Z < 1.14)

= 1 - 0.8729

= 0.1271